Calculate the total number of electrons present in 1.6g of methane

Grade:11

1 Answers

Shalla AB
6 years ago
GMW Of CH4 is 16g Mass is 1.6g No of moles is mass\GMW1.6/16 = 1/10NO of electrons =no of moles × avagadro number1 / 10 x 6.022 x 10 raise to 23that is equal to 6.022 X 10 raise to 22this is the required answer